# TRADES: Generating Realistic Market Simulations with Diffusion Models

### v2

@@ 7.1. Results
Here, we evaluate the usefulness, realism, and the responsiveness of the generated market simulations for Tesla and Intel. We train two TRADES Models, one for each stock. After training, we freeze the models and use them to generate orders during the simulation phase. A major disadvantage of market simulation is the misalignment in the literature for a single evaluation protocol. All the other works (Coletta et al., 2022; Hultin et al., 2023; Nagy et al., 2023; Xia et al., 2024) analyze performances with plots, hindering an objective comparison between the models. To fill this gap we adapt the predictive score (Yoon et al., 2019) for market simulation. Predictive score is measured as an MAE, by training a stock mid-price predictive model on synthetic data and testing it on real data, that evaluates the usefulness of the generated simulations. Appendix D details the computation of the predictive score.
@@
Table 1. Average predictive score (MAE) over two days on Tesla and Intel stocks. Bold values show the best MAE.
@@
Usefulness: TRADES outperforms the second-best by a factor of ×$\mathbf{\times 3.27}$ and ×$\mathbf{\times 3.48}$ on both stocks. [^12] The values are computed dividing the second best value with the TRADES value s, both subtracted by the market replay predictive score. We report both stocks’ average predictive scores in Table 1. We report the predictive score on the market replay as ground truth (market replay). Notice that the market replay scores represent the desired MAE of each model – i.e., the lower the difference in MAE with the market replay, the better. The table reveals that TRADES exhibits performances approaching that of the real market replay, with an absolute difference of $0.29$ and $0.158$ from market replay, respectively, for Tesla and Intel, suggesting a diminishing gap between synthetic and real-data training efficacy. Interestingly, although IABS cannot capture the complexity of real trader agents, it outperforms CGAN on Tesla, while it remains the worst-performing strategy on Intel. Note that the mid-price for Intel is, on average, 1/20th of the mid-price for Tesla. This discrepancy explains the difference in the scale of the predictive score. In conclusion, we demonstrated how a predictive model trained with TRADES’s generated market data can effectively forecast mid-price.
@@
Figure 3. Stylized facts on Tesla 29/01. (1) Log returns autocorrelation. (2) The correlation between volume and volatility, and (3) between returns and volatility. (4) Comparison of the minute Log Returns distribution and (5) autocorrelation. (6) Mid-price traces of five different TRADES simulations.
@@
Figure 4. Average mid-price difference of market replay simulations and TRADES simulations with (shaded part) and without a POV agent (unshaded part), on 5 different seeds.
Figure 5. Volume at the first level of LOB on TSLA 29/01.
@@
Realism: TRADES covers the real data distribution and emulates many stylized facts. To evaluate the realism of the generated time series, we compare the real data distributions with those of the generated one. We employ a combination of Principal Component Analysis (PCA) (Shlens, 2014), alongside specialized financial methodologies that include the comparison of stylized facts and important features. In Fig. 2, we show PCA plots to illustrate how well the synthetic distribution covers the original data. TRADES (blue points) cover 67.04% of the real data distribution (red points) according to a Convex Hull intersection compared to 52.92% and 57.49% of CGAN and IABS, respectively. Following (Vyetrenko et al., 2020), we evaluate the simulation realism by ensuring that the generated market data mimics stylized facts derived from real market scenarios. Fig. 3 (1) illustrates that TRADES, similarly to the market replay and differently to IABS and CGAN, obey the expected absence of autocorrelation, indicating that the linear autocorrelation of asset returns decays rapidly, becoming statistically insignificant after 15 minutes. Fig. 3 (2) highlights TRADES’s resemblance with the positive volume-volatility correlation. TRADES’s generated orders show a positive volume-volatility correlation. Interestingly, TRADES captures this phenomenon better than that particular market replay day. Recall that TRADES is trained on 17 days of the market, which shows this characteristic correctly capturing this correlation. We acknowledge that the real market might not always respect all the stylized facts due to their inherent non-deterministic and non-stationary nature. Also, CGAN resembles this phenomenon but, differently from TRADES, with an unrealistic intensity.
@@
Fig 3 (3) shows that TRADES exhibits asset returns and volatility negative correlation emulating the market replay distribution in contrast with IABS. Similar to the previous case, also, CGAN resembles this phenomenon but, differently from TRADES, with an unrealistic intensity. Fig. 3 (4) illustrates that TRADES almost perfectly resembles the real distribution in terms of log returns. We leave IABS out since it disrupts the plot’s scale. In Fig. 3 (5), we show the autocorrelation function of the squared returns, commonly used as an analytical tool to measure the degree of volatility clustering: i.e., high-volatility episodes exhibit a propensity to occur in close temporal proximity. Note that TRADES emulates the real distribution better than the other two methods. In Fig. 3 (6), we illustrate the mid-price time series of five different TRADES [^13] TRADES is frozen, and the same model is used for all simulations. simulations on the same day of Tesla (29/01) and the market replay of that day. The mid-price traces generated show diversity and realism. Lastly, to consolidate our claims about TRADES’s realism, in Fig. 5 we also analyze the volume distribution of the first LOB level. Notice how the scale and the overall behavior of the volume time series in the market replay strongly correlate with that of TRADES’s simulation, while it is completely different w.r.t. SoTA approaches.
@@
Responsiveness: TRADES is responsive to external agent. The responsiveness of a LOB market simulation generative model is crucial, especially if the objective of the market simulation is to verify the profitability of trading strategies or perform market impact experiments. Generally (Coletta et al., 2021, 2022; Shi and Cartlidge, 2023), the responsiveness of a generator is assessed through a market impact experiment (A/B test) (Cont et al., 2013). Therefore, we conducted an experiment running some simulations w/ and w/o a Percentage-Of-Volume (POV) agent, which wakes up every minute and places a bunch of buy orders, until either $\phi$ shares have been transacted or the temporal window ends. We refer the reader to Appendix F for the details of the settings of this experiment. Fig. 4 depicts the normalized mid-price difference between the simulations w/ and w/o the POV agent for the market replay and TRADES. Results are averaged over 5 runs. As expected, the historical market simulation exhibits only instantaneous impact (Gould et al., 2013), that is the direct effect of the agent’s orders, which rapidly vanishes. Contrarily, the diffusion-based simulations demonstrate substantial deviation from the baseline simulation without the POV agent, altering the price permanently. Quantifying the permanent price impact in real markets poses a significant challenge, as it requires comparing price differences between scenarios where a specific action took place and those where it did not. Such scenario analysis is not feasible with empirical data. However, by using TRADES-generated realistic simulations, this analysis becomes both feasible and measurable. In fact, the simulations allow us to run identical scenarios both with and without additional trader agents which strategy can be fully defined by the user. These types of counterfactual or "what if" experiments can also be used to have an initial analysis of the consequences of changing financial regulations and trading rules, such as price variation limits, short selling regulation, tick size, and usage rate of dark pools. In conclusion, the observed market impact in the TRADES simulations aligns with real market observations (Gould et al., 2013; Bouchaud et al., 2009), enabling the evaluation of trading strategies141414we want to be clear that it technically enables evaluating trading strategies, but it does not assure any profitability in a real market scenario. and counterfactual experiments.
7.2. DDIM sampling
@@
One of the known limitations of diffusion models is the sampling time. Indeed, the generation of a single sample necessitates hundreds of iterative passes through the neural network. In this work, each model was trained using a diffusion process comprising 100 steps. Recently, Denoising Diffusion Implicit Model (DDIM) sampling method was proposed in (Song et al., 2020a) to speed up the generative process. Given that each hour of market simulation required six hours of computation on an RTX 3090, accelerating the simulation process was a relevant improvement. Consequently, we conducted simulations employing DDIM sampling ($\eta=0$), which is deterministic, utilizing a single step for each order. We use the same trained model. The results, presented in Table 2, demonstrate that the performance degradation is significant but not disastrous despite a remarkable 100-fold increase in computational efficiency.
@@
Table 2. Average predictive score (MAE) over two days on Tesla and Intel stocks. DDIM sampling is done with a single step, while DDPM with 100.
@@
Table 3. Predictive score for the ablation (A) and sensitivity (S) analyses for two days of Tesla simulations.
@@ 7.3. Ablation and sensitivity studies
Sensitivity analyses. We aim to verify whether the complexity of the backbone – i.e., the transformer in TRADES – is needed after all. Therefore, we replace the transformer backbone with an LSTM, leaving the augmentation and conditioning invariant. The table shows that the performances degrade by an average of $6.06$ absolute points. This is expected since transformers directly access all other steps in the sequence via self-attention, which theoretically leaves no room for information loss that occurs in LSTMs. Recall that we concatenate the past orders and the LOB snapshots, after the augmentation, into a single tensor and use it to condition the diffusion model – i.e., TRADES (Orig.). We also tried a cross-attention (CA) conditioning strategy – see TRADES w/ CA – between the past orders and the LOB snapshots. TRADES w/ CA reports an average performance loss of $7.814$ absolute points w.r.t. the original architecture. Note that cross-attention limits the model’s capability because the orders cannot attend to each other but only to LOB and vice-versa. Instead, with concatenation and self-attention, every sequence part can attend to every other vector.
8. Conclusion
@@
We proposed the Transformer-based Denoising Diffusion Probabilistic Engine for LOB Simulations (TRADES) to generate realistic order flows conditioned on the current market state. We evaluated TRADES’s realism and responsiveness. We also adapted the predictive score to verify the usefulness of the generated market data by training a prediction model on them and testing it on real ones. This shows that TRADES can cover the real data distribution by 67% on average and outperforms SoTA by ×$\times 3.27$ and ×$\times 3.48$ on Tesla and Intel. Furthermore, our analyses reflect that TRADES correctly abides by many stylized facts used to evaluate the goodness of financial market simulation approaches. We release DeepMarket, a Python framework for market simulation with deep learning and TRADES-LOB, a synthetic LOB dataset composed of TRADES’s generated market simulations. We argue that TRADES-LOB and DeepMarket will have a positive impact on the research community, as almost no LOB data is freely available. We believe that TRADES is a viable market simulation strategy in controlled environments, and further tests must be performed to have a mature evaluation trading strategy protocol.
Appendix A Denoising diffusion probabilistic model
[…]
(11) $q(\mathbf{x}_{0:T}):=q(\mathbf{x}_{0})\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),$
[…]
(12) $\begin{gathered}p_{\theta}(\mathbf{x}_{0:T}):=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}),\\ p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}):=\mathbf{\mathcal{N}}(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t),\mathbf{\Sigma}_{\theta}(\mathbf{x}_{t},t))\end{gathered}$
The reverse process model is trained with the variational lower bound of the likelihood of $\mathbf{x}_{0}$ as in Eq. (13).
(13) $\begin{gathered}\mathcal{L}_{vlb}:=\mathbb{E}_{q}\Bigg{[}\underbrace{-p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{1})}_{\text{$L_{0}$}}\\ +\sum_{t=2}^{T}\underbrace{D_{KL}(q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})\,||\,p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}))}_{\text{$L_{t-1}$}}\\ +\underbrace{D_{KL}(q(\mathbf{x}_{T}|\mathbf{x}_{0})\,||\,p_{\theta}(\mathbf{x}_{T})}_{\text{$L_{T}$}}\Bigg{]}.\end{gathered}$
Since both $q$ and $p_{\theta}$ are Gaussian, $D_{KL}$ (the Kullback–Leibler divergence) can be evaluated in a closed form with only the mean and covariance of the two distributions. Ho et al. (Ho et al., 2020) propose the following reparametrization of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$:
(14) $\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\mkern 2.0mu\overline{\mkern-2.0mu\alpha\mkern-2.0mu}\mkern 2.0mu_{t}}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right),$
where $\mathbf{x}_{t}=\sqrt{\mkern 2.0mu\overline{\mkern-2.0mu\alpha\mkern-2.0mu}\mkern 2.0mu_{t}}\mathbf{x}_{0}\,+\,\sqrt{1-\mkern 2.0mu\overline{\mkern-2.0mu\alpha\mkern-2.0mu}\mkern 2.0mu_{t}}\boldsymbol{\epsilon}$ s.t. $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $\mathbf{\Sigma}_{\theta}(\mathbf{x}_{t},t)=\sigma^{2}_{t}\mathbf{I}$ where
(15) $\sigma^{2}_{t}=\begin{cases}\beta_{1}&t=1\\ \tilde{\beta}_{t}&1<t\leq T\end{cases}\>\>\text{and}\>\tilde{\beta}_{t}=\frac{1-\mkern 2.0mu\overline{\mkern-2.0mu\alpha\mkern-2.0mu}\mkern 2.0mu_{t-1}}{1-\mkern 2.0mu\overline{\mkern-2.0mu\alpha\mkern-2.0mu}\mkern 2.0mu_{t}}\beta_{t},$
where $\boldsymbol{\epsilon}_{\theta}$ is the trainable function approximated by the neural network, intended to predict $\boldsymbol{\epsilon}$ from $\mathbf{x}_{t}$. As shown in (Song and Ermon, 2019), the denoising function given by Eq. (14) is equivalent to a score model rescaled for score-based generative models. Using this parameterization, Ho et al. (Ho et al., 2020) demonstrated that the inverse process can be learned by minimizing the simplified objective function in Eq. (16).
(16) $\mathcal{L}_{\text{simple}}(\theta):=\mathbb{E}_{t,\mathbf{x}_{0},\boldsymbol{\epsilon}}\Bigr{[}||\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)||^{2}\Bigr{]}$
The denoising function $\boldsymbol{\epsilon}_{\theta}$ aims to recover the noise vector $\boldsymbol{\epsilon}$ that corrupted its input $\mathbf{x}_{t}$. This training objective can be interpreted as a weighted variant of denoising score matching, a method for training score-based generative models (Song and Ermon, 2019; Song et al., 2020b).
Appendix B Overview of FI-2010 dataset
The FI-2010 dataset (Ntakaris et al., [n. d.]) is the most used LOB dataset in the field of the deep learning application to limit order book (Zhang et al., 2019, 2020; Tsantekidis et al., 2017a, b), especially for forecasting tasks. It contains LOB data from five Finnish companies listed on the NASDAQ Nordic stock market: i.e., Kesko Oyj, Outokumpu Oyj, Sampo, Rautaruukki, and Wärtsilä Oyj. The data covers 10 trading days from June 1st to June 14th, 2010. It records about 4 million limit order snapshots for 10 levels of the LOB. The authors sample LOB observations every 10 events, totaling 394,337 events. The label of the data, representing the mid-price movement, depends on the percentage change between the actual price $p_{t}$ and the average of the subsequent $h$ (chosen horizon) mid-prices:
$l_{t}=\frac{m_{+}(t)-p_{t}}{p_{t}}$
The labels are then decided based on a threshold ($\theta$) for the percentage change ($l_{t}$). If $l_{t}>\theta$ or $l_{t}<-\theta$, the label is an up in the former case, and a down in the latter. When $-\theta<l_{t}<\theta$, the label is stationary. The dataset provides the time series and the classes for five horizons $h\in H=\{1,2,3,5,10\}$. The authors of the dataset employed a single threshold $\theta=2\times 10^{-3}$ for all horizons, balancing only the case for $h=5$.
Besides the absence of message files, FI-2010 comes already pre-processed such that the original LOB cannot be reconstructed, impeding comprehensive experimentation. Finally, as shown in (Zhang et al., 2019), this method of labeling the data is susceptible to instability.
Appendix C Implementation Details
All the experiments are run with an RTX 3090 and an A100. We implemented early stopping with a patience of 5 epochs and then used the checkpoint with the best validation loss for simulation. On average we trained each model for 70,000 steps until convergence. Lastly, we halve the learning rate each time the validation loss exceeds the previous one.
Pre-processing. To properly train the neural network, we preprocess the data. The data contains discrete and continuous features, so we rely on different preprocessing methods for each feature type. We replace the timestamp with the time distance from the previous order, which we normalize according to the z-score. We also perform a z-score normalization of message and order book file volume and prices. We encode the event type with an embedding layer. Finally, we remove the order ID and add the depth [^15] The depth is the difference between the order price and the current best available price in the LOB. to the orders feature vector.
Post-processing. The diffusion output is post-processed before being handled through the exchange simulation. First, we report the continuous features (offset, size, depth) to the original scale using the mean and standard deviation computed on the training set; the direction is discretized by simply checking if it is $>0$ (buy) or $<0$ (sell). Lastly, the order type is discretized between 0 (limit order), 1 (cancel order), and 2 (market order) based on the index of the nearest [^16] The distance is computed using the Euclidean distance. embedding layer row. If the order is a limit order, the depth is utilized as the price. For instance, if the depth is 10 and the direction is “buy”, the order will be positioned at a 10-cent difference from the best available bid price. Occasionally, it happens that the size is negative or the depth is over the first 10 levels; in that case, the order is discarded, and a new one is generated. Approximately 25% of the time, the generated cancel order does not directly correspond to an active limit order with the same depth. We identify the limit order with the closest depth and size in such instances.
Hyperparameters search. To find the best hyperparameters, we employ a grid search exploring different values as shown in Table 4. Furthermore, we set the number of diffusion steps to 100. Lastly, We set $\lambda=0.01$ to prevent $\mathcal{L}_{\mathbf{\Sigma}}$ from overwhelming $\mathcal{L}_{\mathbf{\epsilon}}$. We implement this mixed training by relying on the stop gradient functionality (Paszke et al., 2019), in such a way that $\mathcal{L}_{\boldsymbol{\epsilon}}$ optimizes the error prediction and $\mathcal{L}_{\mathbf{\Sigma}}$ the standard deviation.
Table 4. The hyperparameter search spaces and best choice.
Hyperparameter Search Space Best Choice
Optimizer {Adam, Lion} Adam
Sequence size {64, 128, 256, 512} 256
Learning Rate $\{{10^{-3}},10^{-4}\}$ $10^{-3}$
TRADES Layers $\{4,6,{8},16\}$ $8$
Dropout $\{0,{0.1}\}$ $0.1$
Attention Heads $\{1,{2},4\}$ $2$
Augmentation Dim. $\{32,{64},128,256\}$ $64$
$\lambda$$\{0.1,0.01,0.001\}$ $0.01$
Noise scheduler. As shown in (Nichol and Dhariwal, 2021), too much noise at the end of the forward noising process lowers the sample’s quality. Hence, we rely on a non-linear noise scheduler as described in Eq. (17).
(17) $\mkern 2.0mu\overline{\mkern-2.0mua_{t}\mkern-2.0mu}\mkern 2.0mu=\frac{f(t)}{f(0)},\>\>f(t)=\cos{\bigg{(}\frac{t/T+s}{1+s}\cdot\frac{\pi}{2}\bigg{)}}^{2}$
Importance Sampling. Because some diffusion steps contribute to most of the loss, we exploit importance sampling (Nichol and Dhariwal, 2021) to focus on these steps as in Eq. (18).
(18) $\mathcal{L}_{\mathbf{\Sigma}}=\mathbf{E}_{t\sim p_{t}}\bigg{[}\frac{\mathcal{L}_{t}}{p_{t}}\bigg{]},\>\>\>\text{where}\>\>p_{t}\propto\sqrt{\mathbf{E}[\mathcal{L}_{t}^{2}]}\>\>\>\text{and}\>\>\sum_{t}p_{t}=1.$
Diffusion step and positional embedding. We embed the diffusion step $t$ and each vector’s position in the sequence using sinusoidal embedding (Vaswani et al., 2017). Obviously, the diffusion step $t$ embedding is one for the whole sequence, while the position embedding is different for each element.
CGAN implementation. We implemented CGAN from scratch given that none of the implementations in (Coletta et al., 2023, 2022, 2021) are available. Furthermore, we performed a hyperparameters search because the majority of the hyperparameters are not specified. The generator comprises an LSTM, a fully connected, and four 1D convolution layers with a final tanh activation function. Each convolution layer is interleaved with batch normalization and ReLUs. The kernel size of each convolution is 4, and the stride is 2. The optimizer, the learning rate, and the sequence size are the same as TRADES.
An important detail is how the discrete features are post-processed after the tanh function during the generation. We set the binomial feature (direction, quantity type) to -1 if the value is less than 0 and vice versa. While, for the order type, we suppose [^17] The original paper lacks necessary details. that the authors search for a threshold that resembles the real order type distribution. Our final strategy for Tesla is: if the value is lower than 0.1, we assign -1 (limit order); if the value is between 0.1 and 0.25, we set it to 0 (cancel order) and with 1 (market order) otherwise. For Intel, we do smaller changes: if the value is lower than 0.15, we assign -1 (limit order); if the value is between 0.15 and 0.95, we set it to 0 (cancel order) and with 1 (market order) otherwise. The distribution of the generated data is similar to the real one when exploiting this heuristic. We did our best to implement CGAN most competitively. The full implementation is available in our framework.
Appendix D Predictive Score Calculation
We rely on the predictive score (Yoon et al., 2019) to measure how much the synthetic data is effective for the stock mid-price forecasting task. It is computed by training a predictive model $\Phi$ on synthetic data and measures the MAE on a real test set. The task considered is forecasting the mid-price with a horizon of 10, given in input the last 100 market observations. A market observation contains the last order and the first level of the LOB. We choose a 2-layered LSTM (Hochreiter and Schmidhuber, 1997), standard architecture for time series forecasting (Lindemann et al., 2021), for $\Phi$. We train a $\Phi$ on the generated market simulation for each generative method and each simulated day (29/1 and 30/01 for both stocks) for 100 epochs at maximum. We used early stopping with 5 patience. Next, we evaluate each $\Phi$ on the real test set extracted from the market replay. [^18] The market replay is the simulation performed with the real orders of that trading day. In addition, a comparative $\Phi$ model was trained and tested exclusively on real market data to benchmark performance differences.
Appendix E Additional Results
For completeness, we show the volume (see Fig. 6) and stylized facts (see Fig. 7) on 30/01. Notice how, as shown in Fig. 5, the volume for 30/01 on Tesla follows realistic trends. Meanwhile, SoTA approaches cannot seem to cope with the GT. Similar reasoning applies to the stylized facts discussed in the main paper (see Fig. 3). Notice how all of them are satisfied, except returns and volatility negative correlation. Interestingly, the correlation between the volume and volatility (2) is slightly negative for the GT. Nevertheless, because TRADES observes a positive correlation during training, this is also reflected in the generated orders.
Figure 6. Volume at 1st LOB level extracted from the simulation of TSLA on 30/01
Figure 7. Stylized facts of Tesla on 30/01. (1) Log returns autocorrelation. (2) The correlation between volume and volatility, and (3) between returns and volatility. (4-5) Comparison of the minute Log Returns distribution and autocorrelation.
Appendix F Responsiveness Experiment Settings
A $\delta$-POV strategy is characterized by a percentage level $\delta\in(0,1]$, a wake-up frequency $\Delta t$, a direction (buy or sell), and a target quantity of shares $\phi$. The agent wakes up every $\Delta t$ time unit and places buy or sell orders for several shares equal to $\delta V_{t}$. This process continues until either $\phi$ shares have been transacted or the temporal window ends. In our experiments, we use $\delta=0.1$ and a wake-up frequency $\Delta t=1$ min. The temporal window is set from 09:45 to 10:30, and we set $\phi=10^{5}$.

### v3

@@ 7.1 Results
Here, we evaluate the usefulness, realism, and the responsiveness of the generated market simulations for Tesla and Intel. We train two TRADES Models, one for each stock. After training, we freeze the models and use them to generate orders during the simulation phase. A major disadvantage of market simulation is the misalignment in the literature for a single evaluation protocol. All the other works [coletta2022learning, hultin2023generative, nagy2023generative, coletta2021towards] analyze performances with plots, hindering an objective comparison between the models. To fill this gap, we adapt the predictive score [yoon2019time] for market simulation. Predictive score is measured as an MAE, by training a stock mid-price predictive model on synthetic data and testing it on real data, that evaluates the usefulness of the generated simulations. In the Appendix [bertiecaiappendix], we detail the computation of the predictive score.
@@
Table 1: Average predictive score (MAE) over two days on Tesla and Intel stocks. Bold values show the best MAE.
@@
Usefulness: TRADES outperforms the second-best by a factor of $\mathbf{\times 3.27}$ and $\mathbf{\times 3.48}$ on both stocks. [^9] The values are computed dividing the second best value with the TRADES values, both subtracted by the market replay predictive score. We report both stocks’ average predictive scores in Table 1. Notice that the market replay scores represent the desired MAE of each model. The table reveals that TRADES exhibits performances approaching that of the real market replay, with an absolute difference of $0.29$ and $0.158$ from market replay, respectively, for Tesla and Intel, suggesting a diminishing gap between synthetic and real-data training efficacy. Interestingly, although IABS cannot capture the complexity of real trader agents, it outperforms CGAN on Tesla, while it remains the worst-performing strategy on Intel. In conclusion, we demonstrated how a predictive model trained with TRADES’s generated market data can effectively forecast mid-price.
@@
Figure 3: Stylized facts on Tesla 29/01. (1) Log returns autocorrelation. (2) The correlation between volume and volatility, and (3) between returns and volatility. (4) Comparison of the minute Log Returns distribution and (5) autocorrelation. (6) Mid-price traces of five different TRADES simulations.
Figure 4: Volume at the first level of LOB on TSLA 29/01.
@@
Figure 5: Average mid-price difference of market replay simulations and TRADES simulations with (shaded part) and without a POV agent (unshaded part), on 5 different seeds.
Realism: TRADES covers the real data distribution and emulates many stylized facts. To evaluate the realism of the generated time series, we compare the real data distributions with those of the generated one. We employ a combination of Principal Component
@@
Analysis (PCA), alongside specialized financial methodologies that include the comparison of stylized facts and important features. In Fig. 2, we show PCA plots to illustrate how well the synthetic distribution covers the original data. TRADES (blue points) cover 67.04% of the real data distribution (red points) according to a Convex Hull intersection compared to 52.92% and 57.49% of CGAN and IABS, respectively. Following [vyetrenko2020get], we evaluate the simulation realism by ensuring that the generated market data mimics stylized facts derived from real market scenarios. Fig. 3 (1) illustrates that TRADES, similarly to the market replay and differently to IABS and CGAN, obey the expected absence of autocorrelation, indicating that the linear autocorrelation of asset returns decays rapidly, becoming statistically insignificant after 15 minutes. Fig. 3 (2) highlights TRADES’s resemblance with the positive volume-volatility correlation. TRADES’s generated orders show a positive volume-volatility correlation. Interestingly, TRADES captures this phenomenon better than that particular market replay day. Recall that TRADES is trained on 17 days of the market, so it correctly captures this correlation. We acknowledge that the real market might not always respect all the stylized facts due to their inherent non-deterministic and non-stationary nature. Also, CGAN resembles this phenomenon but, differently from TRADES, with an unrealistic intensity.
@@
Fig 3 (3) shows that TRADES exhibits asset returns and volatility negative correlation, emulating the market replay distribution in contrast with IABS. Similar to the previous case, also, CGAN also resembles this phenomenon but with an unrealistic intensity. Fig. 3 (4) illustrates that TRADES almost perfectly resembles the real distribution in terms of log returns. We leave IABS out since it disrupts the plot’s scale. In Fig. 3 (5), we show the autocorrelation function of the squared returns, commonly used as an analytical tool to measure the degree of volatility clustering: i.e., high-volatility episodes exhibit a propensity to occur in close temporal proximity. Note that TRADES emulates the real distribution better than the other two methods. In Fig. 3 (6), we illustrate the mid-price time series of five different TRADES [^10] TRADES is frozen, and the same model is used for all simulations. simulations on the same day of Tesla (29/01) and the market replay of that day. The mid-price traces generated show diversity and realism. Lastly, to consolidate our claims about TRADES’s realism, in Fig. 4 we also analyze the volume distribution of the first LOB level. Notice how the scale and the overall behavior of the volume time series in the market replay strongly correlate with that of TRADES’s simulation, while it is completely different w.r.t. SoTA approaches.
@@
Responsiveness: TRADES is responsive to external agent. The responsiveness of a LOB market simulation generative model is crucial, especially if the objective of the market simulation is to verify the profitability of trading strategies or perform market impact experiments. Generally [coletta2021towards, coletta2022learning, shi2023neural], the responsiveness of a generator is assessed through a market impact experiment (A/B test). Therefore, we conducted an experiment running some simulations w/ and w/o a Percentage-Of-Volume (POV) agent, which wakes up every minute and places a bunch of buy orders, until either $\phi$ shares have been transacted or the temporal window ends. We refer the reader to the Appendix [bertiecaiappendix] for the details of the settings of this experiment. Fig. 5 depicts the normalized mid-price difference between the simulations w/ and w/o the POV agent for the market replay and TRADES. Results are averaged over 5 runs. As expected, the historical market simulation exhibits only instantaneous impact [gould2013limit], that is the direct effect of the agent’s orders, which rapidly vanishes. Contrarily, the diffusion-based simulations demonstrate substantial deviation from the baseline simulation without the POV agent, altering the price permanently. Quantifying the permanent price impact in real markets poses a significant challenge, as it requires comparing price differences between scenarios where a specific action took place and those where it did not. Such scenario analysis is not feasible with empirical data. However, by using TRADES-generated realistic simulations, this analysis becomes both feasible and measurable. In fact, the simulations allow us to run identical scenarios both with and without additional trader agents whose strategy can be fully defined by the user. In conclusion, the observed market impact in the TRADES simulations aligns with real market observations [gould2013limit, bouchaud2009markets], enabling the evaluation of trading strategies111111we want to be clear that it technically enables evaluating trading strategies, but it does not assure any profitability in a real market scenario. and counterfactual experiments.
7.2 DDIM sampling
@@
One of the known limitations of diffusion models is the sampling time. Indeed, the generation of a single sample necessitates hundreds of iterative passes through the neural network. In this work, the model was trained using a diffusion process comprising 100 steps. Recently, the Denoising Diffusion Implicit Model (DDIM) sampling method was proposed in [song2020denoising] to speed up the generative process. Given that each hour of market simulation required six hours of computation on an RTX 3090, accelerating the simulation process was a relevant improvement. Consequently, we conducted simulations employing DDIM sampling ($\eta=0$), which is deterministic, utilizing a single step for each order. We use the same trained model. The results, presented in Table 2, demonstrate that the performance degradation is significant but not disastrous despite a remarkable 100-fold increase in computational efficiency.
@@
Table 2: Average predictive score (MAE) over two days on Tesla and Intel stocks. DDIM sampling is done with a single step, while DDPM with 100.
@@
7.3 Ablation and sensitivity studies
Table 3: Predictive score for the ablation (A) and sensitivity (S) analyses for two days of Tesla simulations.
@@ 7.3 Ablation and sensitivity studies
Sensitivity analyses. We aim to verify whether the complexity of the backbone – i.e., the transformer in TRADES – is needed after all. Therefore, we replace the transformer backbone with an LSTM, leaving the augmentation and conditioning invariant. The table shows that the performances degrade by an average of $6.06$ absolute points. This is expected since transformers directly access all other steps in the sequence via self-attention, which theoretically leaves no room for information loss that occurs in LSTMs. Recall that we concatenate the past orders and the LOB snapshots, after the augmentation, into a single tensor and use it to condition the diffusion model – i.e., TRADES (Orig.). We also tried a cross-attention (CA) conditioning strategy, – see TRADES w/ CA – between the past orders and the LOB snapshots. TRADES w/ CA reports an average performance loss of $7.814$ absolute points w.r.t. the original architecture. Note that cross-attention limits the model’s capability because the orders cannot attend to each other but only to LOB and vice-versa.
8 Conclusion
@@
We proposed a Transformer-based Denoising Diffusion Probabilistic Engine for LOB Simulations (TRADES) to generate realistic order flows conditioned on the current market state. We adapted the predictive score to verify the usefulness of the generated market data by training a prediction model on it and testing it on real data. We showed that TRADES can cover the real data distribution by 67% on average and outperforms SoTA by $\times 3.27$ and $\times 3.48$ on Tesla and Intel. Furthermore, we demonstrated that TRADES correctly abides by many stylized facts used to evaluate the goodness of financial market simulation approaches. We release DeepMarket, a Python framework for market simulation with deep learning and TRADES-LOB, a synthetic LOB dataset composed of TRADES’s generated market simulations. We argue that TRADES-LOB and DeepMarket will have a positive impact on the research community, as almost no LOB data is freely available. We believe that TRADES is a viable market simulation strategy in controlled environments, and further tests must be performed to have a mature evaluation trading strategy protocol. A promising experiment could involve training a larger model across various stocks, eliminating the need for an individual model for each stock. {ack} This work was supported by a grant from Lazio Region, FESR Lazio 2021-2027, project @HOME (F89J23001050007 CUP B83C23006240002) and by the project “E-DAI”: Piano Operativo Salute (POS) 2014-2020, CUP: B83C22004150001.